In the morning, Maria drove to a business appointment at 90 mph. Her average speed on the return trip in the afternoon was 80 mph. The return trip took 1/8 hr longer because of heavy traffic. How far did she travel to the appointment?
Found 2 solutions by Boreal, Alan3354:
Answer by Boreal(15235)

distance=rate * time
time to appt=x
time home= x + (1/8) units are hours.
distance equal
so 90*x=80(x+1/8)
90x=80x +10
10x=10
x =1 hour
Distance is 90 miles. One hour at 90 mph, 1 1/8 hours at 80 mph.

Answer by Alan3354(69443)

80 mi/hr in heavy traffic can be interesting.